On May 15, 2019 Boise Inc. (BMS) paid 320m cash for a thallium smelter. The company engineers estimated that the facility would have a 16 year service life and that its retirement in accordance with then-current Idaho environmental regulations would cost 80m. The risk free rate was 2.8% and BMS, a BB- credit, expects a 550 bps credit spread to rF.

On March 1, 2025 BMS sold the smelter for 190m. The buyer paid cash and assumed the ARO.

QUESTION: Book the 2025 03 01 transfer journal entry. Assume that BMS accountants depreciate the smelter SLN with zero salvage, that discounting was calculated on a continuously compounded basis and that the ARO assumptions hd not changed.
Answer #1

Answer :

Cash paid to purchase Thallium Smelter $320.00 m
(+) PV of retirement cost (Note) $21.20 m
Total cost of the asset $341.20 m

Working Note

The appropriate compound interest formula is

Present value = Ae-rt

Where, A = amount paid at the end of tenure = $80 m

Where, r = rate of interest = (2.80 + 5.50)% = 8.30%

Where, t = tenure = 16 years

Using the above formula,

PV = 80m* e-0.083*16

PV = 80m* e-1.328

PV = 80m* 0.265

PV = 21.20m

Computation of Book value of machine on March 1,2025

Total cost $341.20 m
Salvage value 0
Expected useful life (n) (16*12) 192 months
Period in use (from may 15,2019 to Mar 1,2025 69.5 months
Depreciation upto Mar 1,2025 (341.20*69.50/192) $123.51 m
Book value as on Mar 1, 2025 (341.20 - 123.51) $217.69 m
(-) Cash received on sale $-190.00 m
Loss on transfer (217.69 -190) $27.69 m
Journal
Date Particulars

Debit

(in $ m)

Credit

(in $ m)

01 Mar.25 Cash 190.00
Accumulated Depreciation 123.51
Loss on sale of asset 27.69
Thallium Smelter Machine 341.20
(To transfer of smelter machine for cash)